OpenStudy (anonymous):

simplify by removing factors 1 z^2-36/6-z

OpenStudy (anonymous):

=(z-6)(z+6) / 6-z = (z-6)(z+6) / -(z-6) = -(z+6)

OpenStudy (anonymous):

\[\frac{(z-6)(z+6)}{6-z}\]

OpenStudy (anonymous):

\[-\frac{(z+6)(z-6)}{z-6}=-(z+6)\]

OpenStudy (anonymous):

what do you use to get that format cal?